solve the system of linear equations by graphing.
4x+y=0
2x+y=2

Answer by Alan3354(69443)

solve the system of linear equations by graphing.
4x+y=0
2x+y=2
-----------
They're straight lines, so you need 2 points.
4x+y=0
y = -4x
Pick 2 values for x and find y --> points (x,y)
Plot the 2 points, draw a line thru them.
==============
Do the same for the other eqn.
Where they intersect is the solution, the point where x & y fit both equations.